# Establishing a Foundation for Tetun Text Ad-Hoc Retrieval: Stemming, Indexing, Retrieval, and Ranking

### v2

@@ 11. Acknowledgments
This work is financed by National Funds through the Portuguese funding agency, FCT - Fundação para a Ciência e a Tecnologia, under the Ph.D. scholarship grant number SFRH/BD/151437/2021 (DOI 10.54499/SFRH/BD/151437/2021).
@@ 12.1. The Labadain-Stopwords
The Labadain-Stopwords with their English translations is presented in Table 27.
@@ 12.2. Details of the Labadain-Stemmer Algorithms
eza, ezas, iku, ika, ikus, ikas, izmu, izmus, ável, ível, ista, istas, ozu, oza, ozus, ozas, amentu, amentus, imentu, imentus, adora, adór, asaun, adoras, adores, asoens, ante, antes, ánsia, atória, atóriu, atórias, atórius, amentál $general\_suf$ A list contains general suffixes
lojia, lojias $lojia\_suf$ A list contains loj and lojia suffixes
usaun, usoens $usaun\_suf$ A list contains usaun and usoens suffixes
énsia, énsias $ensia\_suf$ A list contains énsia and énsias suffixes
amente $amente\_suf$ A string with amente value
iv (appears before the amente suffix) $iv\_suf$ A string with iv value
at (takes precedence over the iv, ivaivuivas, or ivus suffixes) $at\_suf$ A string with at value
oz, ik, ad (presents before the amente suffix) $ozikad\_suf$ A list contains ozik, and ad suffixes
mente $mente\_suf$ A string with mente value
ante, avel, ivel (appears before the mente suffix) $ante\_suf$ A list contains anteavel, and ivel suffixes
idade, idades $idade\_suf$ A list contains idade and idades suffixes
abil, is, iv (takes precedence over the idade or idades suffixes) $abil\_suf$ A list contains abilis, and iv suffixes
iva, ivu, ivas, ivus $iva\_suf$ A list contains ivaivuivas and ivus suffixes
ada, adu, adas, adus, ida, idu, idas, idus, ária, áriu, árias, árius $verb\_suf$ A list contains verb suffixes
a, e, i, u, us, as $residual\_suf$ A list contains residual suffixes
@@
$R1$, $R2$, $RV$, $word\_list$
2:$general\_suf\leftarrow$ list of general suffixes
3:$lojia\_suf\leftarrow$ list of lojia suffixes
4:$usaun\_suf\leftarrow$ list of usaun suffixes
5:$ensia\_suf\leftarrow$ list of ensia suffixes
6:$amente\_suf\leftarrow$ amente suffix
7:$iv\_suf\leftarrow$ iv suffix
8:$at\_suf\leftarrow$ at suffix
9:$ozikad\_suf\leftarrow$ list of ozikad suffixes
10:$mente\_suf\leftarrow$ mente suffix
11:$ante\_suf\leftarrow$ list of ante suffixes
12:$idade\_suf\leftarrow$ list of idade suffixes
13:$abil\_suf\leftarrow$ list of abil suffixes
14:$iva\_suf\leftarrow$ list of iva suffixes
15:$verb\_suf\leftarrow$ list of verb suffixes
16:$residual\_suf\leftarrow$ list of residual suffixes
17:for all $word$ in $word\_list$ do $\triangleright$ Step 1: Word length validation
@@
19:         Return $word$
20:     $\triangleright$ Step 2: Standard suffix removal
21:         if $word$ ends with any $suffix$ in $general\_suf$ sorted by length descending then
22:              if Position of $suffix$ in $word$ is in $R2$ then
23:                  Delete $suffix$ from $word$
24:                  $stem\leftarrow$ $word$ without $suffix$
25:                  Return $stem$
@@
27:         else if $word$ ends with any $suffix$ in $lojia\_suf$ then
28:              if Position of $suffix$ in $word$ is in $R2$ then
29:                  Replace $suffix$ in $word$ with $loj$
30:                  $stem\leftarrow$ $word$ without $suffix$ concatenates with $loj$
31:                  Return $stem$
@@
33:         else if $word$ ends with any $suffix$ in $usaun\_suf$ then
34:              if Position of $suffix$ in $word$ is in $R2$ then
35:                  Replace $suffix$ in $word$ with $u$
36:                  $stem\leftarrow$ $word$ without $suffix$ concatenates with $u$
37:                  Return $stem$
@@
39:         else if $word$ ends with any $suffix$ in $ensia\_suf$ then
40:              if Position of $suffix$ in $word$ is in $R2$ then
41:                  Replace $suffix$ in $word$ with $ente$
42:                  $stem\leftarrow$ $word$ without $suffix$ concatenates with $ente$
43:                  Return $stem$
@@
45:         else if $word$ ends with $suffix$ equals $amente\_suf$ then
46:              if Position of $suffix$ in $word$ is in $R1$ then
47:                  Delete $suffix$ in $word$
48:                  $stem\_amente\leftarrow$ $word$ without $suffix$
49:                  if $stem\_amente$ preceded by $iv\_suf$ and the position of $iv\_suf$ is in $R2$ then
50:                       Delete $iv\_suf$ in $stem\_amente$
51:                       $stem\_iv\leftarrow$ $stem\_amente$ without $iv\_suf$
52:                       if $stem\_iv$ further proceeded by $at\_suf$ and the position of $at\_suf$ is in $R2$ then
53:                           Delete $at\_suf$ in $stem\_iv$
54:                           $stem\_at\leftarrow$ $stem\_iv$ without $at\_suf$
55:                           Return $stem\_at$
@@
57:                       Return $stem\_iv$
58:                  else if $stem\_amente$ preceded by any $suffix\_oid$ in $ozikad\_suf$ then
59:                       if Position of $suffix\_oid$ in $stem\_amente$ is in $R2$ then
60:                           Delete $suffix\_oid$ in $stem\_amente$
61:                           $stem\_ozikad\leftarrow$ $stem\_amente$ without $suffix\_oid$
62:                           Return $stem\_ozikad$
@@
65:                  Return $stem\_amente$
@@
67:         else if $word$ ends with $suffix$ equals $mente\_suf$ then
68:              if Position of $suffix$ in $word$ is in $R2$ then
69:                  Delete $suffix$ in $word$
70:                  $stem\_mente\leftarrow$ $word$ without $suffix$
71:                  if $stem\_mente$ preceded by any $suffix\_ant$ in $ante\_suf$ then
72:                       if Position of $suffix\_ant$ in $stem\_mente$ is in $R2$ then
73:                           Delete $suffix\_ant$ in $stem\_mente$
74:                           $stem\_ante\leftarrow$ $stem\_mente$ without $suffix\_ant$
75:                           Return $stem\_ante$
@@
78:                  Return $stem\_mente$
@@
80:         else if $word$ ends with $suffix$ equals $idade\_suf$ then
81:              if Position of $suffix$ in $word$ is in $R2$ then
82:                  Delete $suffix$ in $word$
83:                  $stem\_idade\leftarrow$ $word$ without $suffix$
84:                  if $stem\_idade$ preceded by any $suffix\_abl$ in $abil\_suf$ then
85:                       if Position of $suffix\_abl$ in $stem\_idade$ is in $R2$ then
86:                           Delete $suffix\_abl$ in $stem\_idade$
87:                           $stem\_abil\leftarrow$ $stem\_idade$ without $suffix\_abl$
88:                           Return $stem\_abil$
@@
91:                  Return $stem\_idade$
@@
93:         else if $word$ ends with $suffix$ equals $iva\_suf$ then
94:              if Position of $suffix$ in $word$ is in $R2$ then
95:                  Delete $suffix$ in $word$
96:                  $stem\_iva\leftarrow$ $word$ without $suffix$
97:                  if $stem\_iva$ preceded by $at\_suf$ and the position of $at\_suf$ is in $R2$ then
98:                       Delete $at\_suf$ in $stem\_iva$
99:                       $stem\_at\leftarrow$ $stem\_iva$ without $at\_suf$
100:                       Return $stem\_at$
@@
102:                  Return $stem\_iva$
103:              end $\triangleright$ Step 3: Verb suffix removal
104:         else if $word$ ends with $suffix$ equals $verb\_suf$ then
105:              if Position of $suffix$ in $word$ is in $RV$ then
106:                  Delete $suffix$ in $word$
107:                  $stem\_verb\leftarrow$ $word$ without $suffix$
108:                  Return $stem\_verb$
109:              end $\triangleright$ Step 4: Residual suffix removal
110:         else if $word$ ends with $suffix$ equals $residual\_suf$ then
111:              if Position of $suffix$ in $word$ is in $RV$ then
112:                  Delete $suffix$ in $word$
113:                  $stem\_residual\leftarrow$ $word$ without $suffix$
114:                  Return $stem\_residual$

### v6

@@ 11. Acknowledgments
This work is financed by National Funds through the Portuguese funding agency, FCT - Fundação para a Ciência e a Tecnologia, under the Ph.D. studentship grant number SFRH/BD/151437/2021 (DOI: 10.54499/SFRH/BD/151437/2021).
@@ 12.1. The Labadain-Stopwords
The Labadain-Stopwords list (de Jesus and Nunes, 2025c) with their English translations are presented in Table 28.
@@ 12.2. Details of the Labadain-Stemmer Algorithms
eza, ezas, iku, ika, ikus, ikas, izmu, izmus, ável, ível, ista, istas, ozu, oza, ozus, ozas, amentu, amentus, imentu, imentus, adora, adór, asaun, adoras, adores, asoens, ante, antes, ánsia, atória, atóriu, atórias, atórius, amentál general_sufgeneral\_suf A list contains general suffixes
lojia, lojias lojia_suflojia\_suf A list contains loj and lojia suffixes
usaun, usoens usaun_sufusaun\_suf A list contains usaun and usoens suffixes
énsia, énsias ensia_sufensia\_suf A list contains énsia and énsias suffixes
amente amente_sufamente\_suf A string with amente value
iv (appears before the amente suffix) iv_sufiv\_suf A string with iv value
at (takes precedence over the iv, ivaivuivas, or ivus suffixes) at_sufat\_suf A string with at value
oz, ik, ad (presents before the amente suffix) ozikad_sufozikad\_suf A list contains ozik, and ad suffixes
mente mente_sufmente\_suf A string with mente value
ante, avel, ivel (appears before the mente suffix) ante_sufante\_suf A list contains anteavel, and ivel suffixes
idade, idades idade_sufidade\_suf A list contains idade and idades suffixes
abil, is, iv (takes precedence over the idade or idades suffixes) abil_sufabil\_suf A list contains abilis, and iv suffixes
iva, ivu, ivas, ivus iva_sufiva\_suf A list contains ivaivuivas and ivus suffixes
ada, adu, adas, adus, ida, idu, idas, idus, ária, áriu, árias, árius verb_sufverb\_suf A list contains verb suffixes
a, e, i, u, us, as residual_sufresidual\_suf A list contains residual suffixes
@@
1:$R1$, $R2$, RVRV, word_listword\_list
[…]
17:for all wordword in word_listword\_list do $\triangleright$ Step 1: Word length validation
@@
19:   Return wordword
[…]
21:   if wordword ends with any suffixsuffix in general_sufgeneral\_suf sorted by length descending then
22:     if Position of suffixsuffix in wordword is in $R2$ then
23:      Delete suffixsuffix from wordword
24:      $stem\leftarrow$ wordword without suffixsuffix
25:      Return stemstem
@@
27:   else if wordword ends with any suffixsuffix in lojia_suflojia\_suf then
28:     if Position of suffixsuffix in wordword is in $R2$ then
29:      Replace suffixsuffix in wordword with lojloj
30:      $stem\leftarrow$ wordword without suffixsuffix concatenates with lojloj
31:      Return stemstem
@@
33:   else if wordword ends with any suffixsuffix in usaun_sufusaun\_suf then
34:     if Position of suffixsuffix in wordword is in $R2$ then
35:      Replace suffixsuffix in wordword with $u$
36:      $stem\leftarrow$ wordword without suffixsuffix concatenates with $u$
37:      Return stemstem
@@
39:   else if wordword ends with any suffixsuffix in ensia_sufensia\_suf then
40:     if Position of suffixsuffix in wordword is in $R2$ then
41:      Replace suffixsuffix in wordword with enteente
42:      $stem\leftarrow$ wordword without suffixsuffix concatenates with enteente
43:      Return stemstem
@@
45:   else if wordword ends with suffixsuffix equals amente_sufamente\_suf then
46:     if Position of suffixsuffix in wordword is in $R1$ then
47:      Delete suffixsuffix in wordword
48:      $stem\_amente\leftarrow$ wordword without suffixsuffix
49:      if stem_amentestem\_amente preceded by iv_sufiv\_suf and the position of iv_sufiv\_suf is in $R2$ then
50:        Delete iv_sufiv\_suf in stem_amentestem\_amente
51:        $stem\_iv\leftarrow$ stem_amentestem\_amente without iv_sufiv\_suf
52:        if stem_ivstem\_iv further proceeded by at_sufat\_suf and the position of at_sufat\_suf is in $R2$ then
53:         Delete at_sufat\_suf in stem_ivstem\_iv
54:         $stem\_at\leftarrow$ stem_ivstem\_iv without at_sufat\_suf
55:         Return stem_atstem\_at
@@
57:        Return stem_ivstem\_iv
58:      else if stem_amentestem\_amente preceded by any suffix_oidsuffix\_oid in ozikad_sufozikad\_suf then
59:        if Position of suffix_oidsuffix\_oid in stem_amentestem\_amente is in $R2$ then
60:         Delete suffix_oidsuffix\_oid in stem_amentestem\_amente
61:         $stem\_ozikad\leftarrow$ stem_amentestem\_amente without suffix_oidsuffix\_oid
62:         Return stem_ozikadstem\_ozikad
@@
65:      Return stem_amentestem\_amente
@@
67:   else if wordword ends with suffixsuffix equals mente_sufmente\_suf then
68:     if Position of suffixsuffix in wordword is in $R2$ then
69:      Delete suffixsuffix in wordword
70:      $stem\_mente\leftarrow$ wordword without suffixsuffix
71:      if stem_mentestem\_mente preceded by any suffix_antsuffix\_ant in ante_sufante\_suf then
72:        if Position of suffix_antsuffix\_ant in stem_mentestem\_mente is in $R2$ then
73:         Delete suffix_antsuffix\_ant in stem_mentestem\_mente
74:         $stem\_ante\leftarrow$ stem_mentestem\_mente without suffix_antsuffix\_ant
75:         Return stem_antestem\_ante
@@
78:      Return stem_mentestem\_mente
@@
80:   else if wordword ends with suffixsuffix equals idade_sufidade\_suf then
81:     if Position of suffixsuffix in wordword is in $R2$ then
82:      Delete suffixsuffix in wordword
83:      $stem\_idade\leftarrow$ wordword without suffixsuffix
84:      if stem_idadestem\_idade preceded by any suffix_ablsuffix\_abl in abil_sufabil\_suf then
85:        if Position of suffix_ablsuffix\_abl in stem_idadestem\_idade is in $R2$ then
86:         Delete suffix_ablsuffix\_abl in stem_idadestem\_idade
87:         $stem\_abil\leftarrow$ stem_idadestem\_idade without suffix_ablsuffix\_abl
88:         Return stem_abilstem\_abil
@@
91:      Return stem_idadestem\_idade
@@
93:   else if wordword ends with suffixsuffix equals iva_sufiva\_suf then
94:     if Position of suffixsuffix in wordword is in $R2$ then
95:      Delete suffixsuffix in wordword
96:      $stem\_iva\leftarrow$ wordword without suffixsuffix
97:      if stem_ivastem\_iva preceded by at_sufat\_suf and the position of at_sufat\_suf is in $R2$ then
98:        Delete at_sufat\_suf in stem_ivastem\_iva
99:        $stem\_at\leftarrow$ stem_ivastem\_iva without at_sufat\_suf
100:        Return stem_atstem\_at
@@
102:      Return stem_ivastem\_iva
[…]
104:   else if wordword ends with suffixsuffix equals verb_sufverb\_suf then
105:     if Position of suffixsuffix in wordword is in RVRV then
106:      Delete suffixsuffix in wordword
107:      $stem\_verb\leftarrow$ wordword without suffixsuffix
108:      Return stem_verbstem\_verb
[…]
110:   else if wordword ends with suffixsuffix equals residual_sufresidual\_suf then
111:     if Position of suffixsuffix in wordword is in RVRV then
112:      Delete suffixsuffix in wordword
113:      $stem\_residual\leftarrow$ wordword without suffixsuffix
114:      Return stem_residualstem\_residual